We roll a pair of dice three times. Find the probability that a total of five is rolled each time. The probability that a total of five is rolled each time is
Answer by stanbon(75887)

We roll a pair of dice three times. Find the probability that a total of five is rolled each time. The probability that a total of five is rolled each time is
------
Pairs:: 1/4 ; 2/3 ; 3/2 ; 4/1
P(5 in one roll) = 4/36 = 1/9
-----
P(5 rolled 3 consecutive times) = (1/9)^3 = 0.00137
------------
Cheers,
Stan H.
-------------